# Tropical Gradient Descent

### v2

@@ 4.2.1 Centrality Statistics
Figure 5: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the Fermat–Weber objective. Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
@@
Table 1: The mean log relative error of each gradient method after 1000 steps for each dataset when minimizing the Fermat–Weber objective function. The minimal mean errors for each dataset are in bold.
@@
6$\times$10 Branching Data -4.60 -4.60 -4.50 -3.32 -4.60 -4.60 -4.60
6$\times$10 Coalescent Data -4.58 -4.55 -4.40 -1.28 -4.57 -4.57 -4.57
6$\times$10 Gaussian Data -4.60 -4.60 -4.44 -2.75 -4.60 -4.60 -4.56
6$\times$100 Branching Data -4.60 -4.60 -4.39 -4.02 -4.54 -4.58 -4.59
6$\times$100 Coalescent Data -4.55 -4.54 -4.25 -2.88 -4.56 -4.58 -4.55
6$\times$100 Gaussian Data -4.60 -4.60 -4.39 -4.03 -4.52 -4.57 -4.59
28$\times$10 Branching Data -4.60 -4.60 -4.15 -1.89 -4.60 -4.60 -4.56
28$\times$10 Coalescent Data -4.53 -3.76 -3.96 -2.11 -4.52 -4.57 -3.87
28$\times$10 Gaussian Data -4.58 -4.58 -4.07 -2.40 -4.60 -4.60 -4.47
28$\times$100 Branching Data -4.59 -4.57 -4.04 -2.89 -4.59 -4.59 -4.59
28$\times$100 Coalescent Data -4.51 -4.54 -3.91 -2.33 -4.39 -4.58 -4.39
28$\times$100 Gaussian Data -4.59 -4.53 -4.10 -2.78 -4.59 -4.59 -4.59
4.2.2 Wasserstein Projections
The Wasserstein projection problems highlight the differences in convergent behavior between classical and tropical gradient methods, as they are $\min$-tropical location problems but not classically convex. The CDFs of the final log relative errors can be found in Figures 6 and 7.
Figure 6: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the 2-Wasserstein projection objective. Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
Figure 7: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the $\infty$-Wasserstein projection objective. Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
In Figures 6 and 7 we see large steps at various intervals in the CDF plots, which is indicative of stable local minima. This is particularly prevalent in low dimensional space. Classical descent, Adam, and Adamax are particularly susceptible to this, while tropical descent and TrAdamax are more likely to pass these local minima. This difference is most pronounced when solving the $\infty$-Wasserstein problem in low dimensions—we are almost certain to reach the minimum with small error, while classical methods have a 50-80% chance of achieving the same error.
4.2.3 Linear Regression
As the coalescent data lies on a tropical hyperplane (Ardila & Klivans,, 2006; Page et al.,, 2020), the minimum of the linear regression function for such data is exactly 0. Figure 8 therefore shows the CDF of the log error rather than the CDF of the log relative error.
@@ 4.2.3 Linear Regression
Figure 8: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the linear regression objective. Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
Table 2: The mean log relative error of each gradient method after 1000 steps for each dataset when minimizing the linear regression objective function. The minimal mean errors for each dataset are in bold.
Data CD TD SGD TSGD Adam Adamax TrAdamax
6$\times$10 Branching Data -0.97 -5.79 -0.56 -2.16 -0.80 -0.84 -4.39
6$\times$10 Coalescent Data -2.70 -5.87 -2.11 -5.47 -2.97 -2.66 -6.58
6$\times$10 Gaussian Data -1.19 -5.48 -0.80 -1.35 -1.05 -1.10 -4.10
6$\times$100 Branching Data -0.79 -5.49 -0.71 -1.43 -0.72 -0.72 -3.53
6$\times$100 Coalescent Data -5.54 -5.76 -3.11 -5.57 -6.90 -7.01 -6.52
6$\times$100 Gaussian Data -1.04 -5.50 -0.88 -1.70 -1.02 -1.02 -3.81
28$\times$10 Branching Data -1.49 -5.51 -0.93 -4.49 -1.14 -1.14 -5.57
28$\times$10 Coalescent Data -2.43 -6.63 -1.77 -5.55 -2.16 -2.29 -6.18
28$\times$10 Gaussian Data -1.26 -5.16 -0.79 -4.29 -1.08 -1.08 -5.72
28$\times$100 Branching Data -0.90 -4.37 -0.61 -1.42 -0.85 -0.77 -3.32
28$\times$100 Coalescent Data -3.33 -5.84 -2.17 -5.35 -3.02 -3.20 -4.93
28$\times$100 Gaussian Data -0.91 -4.35 -0.62 -1.07 -0.96 -0.90 -3.51
@@ 5 Conclusion and Discussion
We have introduced the concept of tropical descent as the steepest descent method with respect to the tropical norm. We have demonstrated, both theoretically and experimentally, that it possesses well-behaved properties with respect to a wide class of optimization problems on the tropical projective torus, including several key statistical problems for the analysis of phylogenetic data. As well as proposing our own framework for tropical gradient methods, this work also provides a detailed review of the varying behavior of popular gradient methods when applied to tropical statistical optimization problems.
@@
Our theoretical results outline the relative strengths of tropical descent over classical descent for tropical location problems, but only provide global solvability for 1-sample problems. In practice, our experimental results go further; the optimization problems studied are location problems, rather than $\triangle$-star-quasi-convex problems, and yet tropical descent appears to consistently converge to a global minimum, not just the tropical convex hull of our dataset. It would therefore be a natural extension of this work to identify how much we can relax the $\triangle$-star-quasi-convexity condition while guaranteeing convergence.
@@
In this paper we have considered tree data with 4 to 8 leaves, and samples of size 10 to 100—at this scale, we have observed (T/)SGD to be a significant sacrifice in accuracy with little benefit in computational time. Furthermore, at this scale, the relative merits of SGD versus its tropical counterpart are inconsistent; TSGD out-performs SGD when solving the linear regression problem, but performs worse when finding centrality statistics. In a similar vein, initial investigations suggest that the classical stochastic algorithm for the computation of Fréchet means—Sturm’s algorithm (Sturm,, 2003)—performs poorly for the computation of tropical Fréchet means due to the inconsistent curvature of the tropical projective torus (Améndola & Monod,, 2023). However, for applications to larger datasets and more complex objective functions, stochastic gradient methods may become necessary. It would then be important to understand the theoretical differences between SGD and TSGD in terms of their convergence guarantees in the tropical setting and their comparative performance at different scales of data dimension.
@@
The recent identification of tropical rational functions and ReLU neural networks (Zhang et al.,, 2018) has inspired new avenues of research in tropical applications, into the tropical geometry of neural networks (Brandenburg et al.,, 2024; Yoshida et al.,, 2024). It is therefore natural to consider the possible extensions of this work to neural network training; can we identify local tropical quasi-convexity in the loss functions of ReLU neural networks, and how can we extend tropical steepest descent to the space of tropical rational functions?
@@
Finally, as shown by (Akian et al.,, 2023), the tropical linear regression problem is polynomial-time equivalent to mean payoff games—two-player perfect information games on a directed graph (Ehrenfeucht & Mycielski,, 1979; Gurvich et al.,, 1988). This family of problems is of particular interest in game theory as they are known to be NP $\cap$ co-NP. In our numerical experiments, tropical gradient methods demonstrate a particularly strong performance in solving tropical linear regression, and while our methods are designed to approximate the minimum rather than compute it exactly, tropical descent may prove to be an efficient tool when solving large-scale mean payoff games.
@@ Acknowledgements
R.T. receives partial funding from a Technical University of Munich–Imperial College London Joint Academy of Doctoral Studies (JADS) award (2021 cohort, PIs Drton/Monod). A.M. is supported by the Engineering and Physical Sciences Research Council under grant reference [EP/Y028872/1] and a London Mathematical Society Emmy Noether Fellowship [EN-2223-01].
@@ Appendix A PCA and Logistic Regression
Figure 9: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the PCA objective. Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
Figure 10: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the logistic regression objective. Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
As well as the location problems outlined in this paper, we performed the same numerical experiments with the tropical polytope PCA loss function (Yoshida et al.,, 2019) and tropical logistic regression loss function (Aliatimis et al.,, 2024). These loss functions take multiple points in $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ as argument, and we execute our gradient methods on each argument simultaneously. Figures 9 and 10 show the resulting CDF of our gradient methods for these problems. These problems do not lend themselves well to gradient methods, as the relative log error is rather large for all of our proposed methods. However, for both problems the tropical descent and TrAdamax methods are outperforming the classical methods quite consistently.
@@ Appendix B Methodology
The same 12 datasets are used for each experiment in Section 4. These datasets live in $\mathbb{R}^{6}/\mathbb{R}\mathbf{1}$ or $\mathbb{R}^{28}/\mathbb{R}\mathbf{1}$; these dimensions correspond to the spaces of trees with 4 and 8 leaves respectively. Branching process datasets are generated using the R package ape (Paradis & Schliep,, 2019) with exponentially distributed branch lengths—this produces data supported on the space of metric trees, i.e., the tropical Grassmannian. Coalescent data has also been generated using ape and is supported on the space of ultrametric trees, which is a tropical linear space. Gaussian data is generated using NumPy (Harris et al.,, 2020), and is supported on the general tropical projective torus.
@@
We use PyTorch (Paszke et al.,, 2019) for the computation and differentiation of each objective function. Max-tropical descent directions are used for tropical gradient methods; we therefore optimize the min-tropical location problems—Wasserstein projections and tropical linear regression—with respect to $-\mathbf{t}$.

### v3

@@ 4.2.2 Wasserstein Projections
Figure 7: The CDF of the log relative error after 1000 steps for each gradient method with tuned learning rates across 50 random (Gaussian) initializations when minimizing the 2-Wasserstein projection objective. Each subfigure corresponds to a dataset with sample size $K=10,100$ of dimensionality $N=6,28$, sampled from a branching process, coalescent process, or a Gaussian distribution.
Figure 8: The CDF of the log relative error after 1000 steps for each gradient method with tuned learning rates across 50 random (Gaussian) initializations when minimizing the $\infty$-Wasserstein projection objective. Each subfigure corresponds to a dataset with sample size $K=10,100$ of dimensionality $N=6,28$, sampled from a branching process, coalescent process, or a Gaussian distribution.
In Figures 7 and 8 we see large steps at various intervals in the CDF plots, which is indicative of stable local minima. This is particularly prevalent in low dimensional space. Classical descent, Adam, and Adamax are particularly susceptible to this, while tropical descent and TrAdamax are more likely to pass these local minima. This difference is most pronounced when solving the $\infty$-Wasserstein problem in low dimensions—we are almost certain to reach the minimum with relative error $\leq e^{-3}$, while classical methods have a 50-80% chance of achieving the same error.
4.2.3 Linear Regression
As the coalescent data lies on a tropical hyperplane (ardila2006bergman; page2020tropical), the minimum of the linear regression function for such data is exactly 0. Figure 9 therefore shows the CDF of the log error rather than the CDF of the log relative error.
@@ 4.2.3 Linear Regression
Table 2: The mean log relative error of each gradient method after 1000 steps with tuned learning rates for branching process, coalescent process, and Gaussian datasets of size $K=10,100$ in $\mathbb{R}^{6}/\mathbb{R}\mathbf{1}$, $\mathbb{R}^{28}/\mathbb{R}\mathbf{1}$ when minimizing the linear regression objective function. The minimal mean errors for each dataset are in bold.
@@
[…]
@@
Figure 9: The CDF of the log error after 1000 steps for each gradient method with tuned learning rates across 50 random (Gaussian) initializations when minimizing the linear regression objective. Each subfigure corresponds to a dataset with sample size $K=10,100$ of dimensionality $N=6,28$, sampled from a branching process, coalescent process, or a Gaussian distribution.
4.3 Tropical Statistics for Applications
We conclude our computational experiments by using tropical gradient methods to solve tropical statistical problems which appear in phylogenetic data analysis and game theory. In particular we look to use tropical centrality statistics to estimate the species tree of a multi-species coalescent model (liu2019modern) , and we use tropical linear regression to estimate the hidden preference factors in the Auction model introduced by akian2023tropical, which is a variant of the first-price sealed-bid auction model (krishna2009auction).
4.3.1 Multi-species Coalescent Model
A species tree refers to a rooted metric tree which represents the evolutionary relationships between a set of species. However, for a specific gene, we may not observe the same branching pattern, instead observing a gene-specific evolution which is referred to as a gene tree. This raises one of the key questions in phylogenetics; how do we estimate the species tree from a set of gene trees? The multi-species coalescent model is a probabilistic model for gene trees given a certain species tree (see liu2019modern for a detailed introduction), which is particularly valuable for formulating Bayesian estimators for species trees and producing simulated data with which to test species tree estimation methodologies.
In this section, we consider the tropical Fermat-Weber and Fréchet means of gene trees as estimators for the underlying species tree. We fix a species tree with 8 leaves (see Figure 10), then use Dendropy (dendropy5) to sample 100 gene trees from a multi-species coalescent model with this underlying species tree and normalize this dataset to have average tropical norm 1.
Figure 10: The underlying species tree used for the multi-species coalescent model generating our dataset of gene trees. The topology of this tree can be represented by the Newick string “(((((F,E),(H,G)),(D,C)),B),A)”.
We then compute a Fermat-Weber point and Fréchet mean for this dataset using each of our gradient methods for 1000 steps with their trained learning rates (Appendix C). We perform these computations from 100 random initializations for each gradient method, then project the result to ultrametric space using the single-linkage algorithm, which is known to coincide with tropical projection (single-linkage; ardila2005subdominant), and record the topology of the resulting tree. The counts for each observed tree topology are presented in Tables 3 and 4.
Newick Tree Adam Adamax CD SGD TD TrAdamax TSGD
((B,A),(((F,E),(H,G)),(D,C))); 0 0 0 0 0 0 2
(((((F,E),(H,G)),(D,C)),A),B); 0 0 0 3 0 0 0
(((((F,E),(H,G)),(D,C)),B),A); 100 100 100 97 100 100 98
Table 3: The counts of each Newick string (up to tree topology) returned by the tropical projection of a Fermat–Weber point computed via each gradient method. Each method is run for 1000 steps with trained learning rate, from 100 random initializations.
Newick Tree Adam Adamax CD SGD TD TrAdamax TSGD
((B,A),(((F,E),(H,G)),(D,C))); 0 0 0 0 0 0 5
(((((F,E),(H,G)),(D,C)),A),B); 0 0 0 1 0 0 1
(((((F,E),(H,G)),(D,C)),B),A); 100 100 100 99 100 100 94
Table 4: The counts of each Newick string (up to tree topology) returned by the tropical projection of a Fréchet means computed via each gradient method. Each method is run for 1000 steps with trained learning rate, from 100 random initializations.
We first note that in the vast majority of cases, our calculations recover the tree topology of the original species tree, suggesting that for this multi-species coalescent dataset, the tropical Fermat–Weber and Frécht mean points are sufficiently good estimators for the species tree. When comparing the performance of each gradient method, we notice that the stochastic methods are the only methods to not recover the species tree topology, with Fréchet means computed via TSGD producing the highest error rate.
The relative performance of tropical Fermat–Weber and Fréchet mean points as estimators for multi-species coalescent species trees is currently an open topic of study (cox2025tropical). In exploring the consistency of these estimators it world be natural for future work to also consider the relative performance of tropical gradient methods for their computation.
4.3.2 Auction Model with Hidden Preference Factors
As a final computational demonstration, we use tropical gradient methods to solve the problem of tropical linear regression using the toy auction data example given by akian2023tropical. For $N$ firms and $K$ products, we let $(p_{ij})_{i\leq N,j\leq K}$ denote the price firm $i$ offers for product $j$. Assuming the offered prices are in a state of equilibrium, that is, no firm can raise their price and still win, differences in proposed price from one firm relative to another is indicative of a hidden preference factor $f_{i}$ in the market; that is, two firms’ offers for product $j$ are considered equal when $p_{i_{1}j}f_{i_{1}}=p_{i_{2}j}f_{i_{2}}$. By performing tropical linear regression on the data $V_{ij}=\exp(-p_{ij})$, we can detect these hidden preference factors. For further details of this model and its assumptions, see akian2023tropical.
For our computations, we take $N=3$ and $K=6,100$, using the same toy datasets discussed in Section 7.2 of akian2023tropical, which have true hidden preference factors $f=(1,0.8,0.6)$. We then perform each gradient method for 1000 steps over 100 random initializations, using learning rates which have been tuned to these datasets (Appendix C). Our estimates for $f$ are then given by $\hat{f}\coloneqq\exp(\mathbf{t}_{1000}-\max(\mathbf{t}_{1000}))$, such that $\max_{i}\hat{f}=1$.
K=6 K=100
Adam (0.76 $\pm$ 0.33, 0.70 $\pm$ 0.32, 0.56 $\pm$ 0.21) (0.77 $\pm$ 0.33, 0.71 $\pm$ 0.32, 0.58 $\pm$ 0.21)
Adamax (0.76 $\pm$ 0.33, 0.70 $\pm$ 0.32, 0.56 $\pm$ 0.21) (0.77 $\pm$ 0.33, 0.71 $\pm$ 0.32, 0.59 $\pm$ 0.21)
CD (0.71 $\pm$ 0.43, 0.60 $\pm$ 0.42, 0.45 $\pm$ 0.30) (0.78 $\pm$ 0.33, 0.70 $\pm$ 0.32, 0.58 $\pm$ 0.20)
SGD (0.73 $\pm$ 0.44, 0.57 $\pm$ 0.43, 0.41 $\pm$ 0.32) (0.77 $\pm$ 0.33, 0.70 $\pm$ 0.32, 0.58 $\pm$ 0.21)
TD (1.00 $\pm$ 0.00, 0.81 $\pm$ 0.02, 0.60 $\pm$ 0.01) (1.00 $\pm$ 0.02, 0.75 $\pm$ 0.08, 0.62 $\pm$ 0.05)
TrAdamax (1.00 $\pm$ 0.00, 0.81 $\pm$ 0.00, 0.60 $\pm$ 0.00) (0.93 $\pm$ 0.19, 0.79 $\pm$ 0.16, 0.65 $\pm$ 0.08)
TSGD (1.00 $\pm$ 0.00, 0.82 $\pm$ 0.02, 0.60 $\pm$ 0.01) (1.00 $\pm$ 0.04, 0.79 $\pm$ 0.04, 0.61 $\pm$ 0.01)
Table 5: The mean($\pm$ 1 std.) estimators $\hat{f}\coloneqq\exp(\mathbf{t}_{1000}-\max(\mathbf{t}_{1000}))$ for the hidden preference factors in the Auction model (akian2023tropical) when computed using gradient methods to solve the tropical linear regression problem. The true hidden preference factors are $f=(1,0.8,0.6)$. akian2023tropical uses the projective Krasnoselskii-Mann iteration to solve this tropical linear regression problem, which produced an estimate of $f^{\text{reg}}=(1,0.81,0.605)$ for the $K=6$ toy dataset.
Table 5 shows the resulting mean estimates for these hidden preference factors when using different gradient methods to solve the tropical linear regression problem. We see that the accuracy of the tropical gradient methods vastly exceeds each of the classical gradient methods, and that these gradient methods produce particularly consistent results for the small sample problem.
These computations are limited to a single Auction model, but show promising results. For future work, it would be of interest to explore how the precision of these methods changes with sample size and dimension, as well as a running time comparison with existing methods such as the projective Krasnoselskii-Mann iteration used by akian2023tropical or combinatorial simplex algorithms (allamigeon2014combinatorial).
@@ 5 Conclusion and Discussion
We have introduced the concept of tropical descent as the steepest descent method with respect to the tropical norm. We have demonstrated, both theoretically and experimentally, that it exhibits desirable properties with respect to a wide class of optimization problems on the tropical projective torus, including several key statistical problems for the analysis of phylogenetic data. As well as proposing our own framework for tropical gradient methods, this work also provides a detailed review of the varying behavior of popular gradient methods when applied to tropical statistical optimization problems.
@@
Our theoretical results outline the relative strengths of tropical descent over classical descent for tropical location problems, but only provide global solvability for 1-sample problems. In practice, our experimental results go much further; the optimization problems studied are location problems, rather than $\triangle_{\min}$-star-quasi-convex problems, and yet tropical descent appears to consistently converge to a global minimum, not just the tropical convex hull of our dataset. It would therefore be a natural extension of this work to identify how much we can relax the $\triangle_{\min}$-star-quasi-convexity condition while guaranteeing global solvability.
@@
In this paper we have considered tree data with 4 to 8 leaves, and samples of size 10 to 100—at this scale, we have observed (T/)SGD to be a significant sacrifice in accuracy with little benefit in computational time. Furthermore, at this scale, the relative merits of SGD versus its tropical counterpart are inconsistent; TSGD out-performs SGD when solving the linear regression problem, but performs worse when finding centrality statistics. In a similar vein, initial investigations suggest that the classical stochastic algorithm for the computation of Fréchet means—Sturm’s algorithm (sturm2003probability)—performs poorly for the computation of tropical Fréchet means due to the inconsistent curvature of the tropical projective torus (amendola2021invitation). However, for applications to larger datasets and more complex objective functions, stochastic gradient methods may become necessary. It would then be important to understand the theoretical differences between SGD and TSGD in terms of their convergence guarantees in the tropical setting and their comparative performance at different scales of data dimension.
@@
The recent identification of tropical rational functions and ReLU neural networks (zhang2018tropical) has inspired new avenues of research in tropical applications, such as the tropical geometry of neural networks (brandenburg2024real; yoshida2023tropNNs). It is therefore natural to consider the possible extensions of this work to neural network training; can we identify local tropical quasi-convexity in the loss functions of ReLU neural networks, and how can we extend tropical descent to the space of tropical rational functions?
@@
Finally, as shown by akian2023tropical, the tropical linear regression problem is polynomial-time equivalent to mean payoff games—two-player perfect information games on a directed graph (ehrenfeucht1979positional; gurvich1988cyclic). This family of problems is of particular interest in game theory as they are known to be NP $\cap$ co-NP. In our numerical experiments, tropical gradient methods demonstrate a particularly strong performance in solving tropical linear regression, and while our methods are designed to approximate the minimum rather than compute it exactly, tropical descent may prove to be an efficient tool when solving large-scale mean payoff games.
@@ Acknowledgements
R.T. receives partial funding from a Technical University of Munich–Imperial College London Joint Academy of Doctoral Studies (JADS) award (2021 cohort, PIs Drton/Monod). A.M. is supported by EPSRC AI Hub on Mathematical Foundations of Intelligence: An “Erlangen Programme” for AI No. EP/Y028872/1.
@@ Appendix A PCA and Logistic Regression
As well as the location problems outlined in this paper, we performed the same numerical experiments with the tropical polytope PCA loss function (yoshida2019tropicalPCA) and tropical logistic regression loss function (aliatimis2023logistic). These loss functions take multiple points in $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ as argument, and we execute our gradient methods on each argument simultaneously. Figures 11 and 12 show the resulting CDF of our gradient methods for these problems. These problems do not lend themselves well to gradient methods, as the relative log error is rather large for all of our proposed methods. However, for both problems the tropical descent and TrAdamax methods are outperforming the classical methods quite consistently.
Figure 11: The CDF of the log relative error after 1000 steps for each gradient method with tuned learning rates across 50 random (Gaussian) initializations when minimizing the PCA objective. Each subfigure corresponds to a dataset with sample size $K=10,100$ of dimensionality $N=6,28$, sampled from a branching process, coalescent process, or a Gaussian distribution.
Figure 12: The CDF of the log relative error after 1000 steps for each gradient method with tuned learning rates across 50 random (Gaussian) initializations when minimizing the logistic regression objective. Each subfigure corresponds to a dataset with sample size $K=10,100$ of dimensionality $N=6,28$, sampled from a branching process, coalescent process, or a Gaussian distribution.
@@ Appendix B Methodology
The same 12 datasets are used for each experiment in Subsections 3.2 and 4. These datasets live in $\mathbb{R}^{6}/\mathbb{R}\mathbf{1}$ or $\mathbb{R}^{28}/\mathbb{R}\mathbf{1}$; these dimensions correspond to the spaces of trees with 4 and 8 leaves respectively. Branching process datasets are generated using the R package ape (paradis2019ape) with exponentially distributed branch lengths—this produces data supported on the space of metric trees, i.e., the tropical Grassmannian. Coalescent data has also been generated using ape and is supported on the space of ultrametric trees, which is a tropical linear space. Gaussian data is generated using NumPy (harris2020array), and is supported on the general tropical projective torus.
@@
We use PyTorch (paszke2019pytorch) for the computation and differentiation of each objective function. Max-tropical descent directions are used for tropical gradient methods; we therefore optimize the min-tropical location problems—Wasserstein projections and tropical linear regression—with respect to $\mathbf{t}$.